One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by mszlmb(115)

One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number
x is bigger, y is smaller
x=12+y
y+2x=39
take the first equation, x=12+y
-12 on both sides: x-12=y
now replace y in the second equation with its equivalent, x-12
x-12+2x=39 +12 both sides
3x=51
x=51/3=17.
X, the larger number, =17.